OpenStudy (anonymous):

HELP onsider the following function. f(t) = 2t2 − 1 Find the average rate of change of the function below over the interval [5, 5.1]. Compare this average rate of change with the instantaneous rates of change at the endpoints of the interval. (at t = 5) (at t = 5.1)

OpenStudy (whpalmer4):

average rate of change from \(a\) to \(b\) is \[\frac{f(b)-f(a)}{b-a}\]

OpenStudy (anonymous):

its to hard for me :/

OpenStudy (whpalmer4):

What's hard? you can evaluate \(2t^2-1\) for various values of \(t\), can't you?

OpenStudy (whpalmer4):

\[f(t) = 2t^2-1\]\[f(5) = 2(5)^2-1 = 2(25)-1 = 50-1=49\] \[f(5.1) = \]?

OpenStudy (anonymous):

f(5.1)= 51.02

OpenStudy (whpalmer4):

Good! Now we know \(f(5) = 49\) and \(f(5.1) = 51.02\)We can find the average rate of change from t = 5 to t = 5.1 by my formula: \[m = \frac{f(b)-f(a)}{b-a}\]where \(b = 5.1,\, b=5\) What is the value of \(m\)?